# Towards Safe and Reliable Autonomous Driving: Dynamic Occupancy Set Prediction

### v1

@@ Abstract
In the rapidly evolving field of autonomous driving, accurate trajectory prediction is pivotal for vehicular safety. However, trajectory predictions often deviate from actual paths, particularly in complex and challenging environments, leading to significant errors. To address this issue, our study introduces a novel method for Dynamic Occupancy Set (DOS) prediction, enhancing trajectory prediction capabilities. This method effectively combines advanced trajectory prediction networks with a DOS prediction module, overcoming the shortcomings of existing models. It provides a comprehensive and adaptable framework for predicting the potential occupancy sets of traffic participants. The main contributions of this research include: 1) A novel DOS prediction model tailored for complex scenarios, augmenting traditional trajectory prediction; 2) The development of unique DOS representations and evaluation metrics; 3) Extensive validation through experiments, demonstrating enhanced performance and adaptability. This research contributes to the advancement of safer and more efficient intelligent vehicle and transportation systems.
@@ I Introduction
The rapid progression of autonomous driving technology is heralding a new era in transportation systems, bringing forth notable enhancements in traffic efficiency and vehicular safety. Yet, ensuring the safety and reliability of autonomous vehicles in the multifaceted real-world scenarios presents a significant challenge. This complexity arises not only from the advanced nature of the technology but also from the necessity to sustain system reliability in intricate environments. The realm of trajectory prediction networks has undergone substantial development, becoming fundamental in boosting the predictive capabilities essential for autonomous vehicle navigation. These networks employ advanced algorithms and extensive datasets, enabling predictions of the movements of traffic participants by considering various factors like vehicle dynamics, road conditions, and traffic patterns. Enhanced prediction accuracy and adaptability of these networks are particularly valuable in complex urban environments or situations involving unpredictable human interactions. However, the complexity of real-world conditions and inherent limitations of predictive models often lead to discrepancies between predicted and actual trajectories, posing a detriment to the reliable assurance of autonomous driving safety.
@@
In response to these challenges, our study introduces an innovative approach for estimating Dynamic Occupancy Sets (DOS) of traffic participants. DOS, which mathematically defines all potential positions a vehicle may occupy within a specified timeframe, plays a pivotal role in reinforcing the safety framework of autonomous driving systems. Traditional set-based prediction methods, often reliant on simplistic dynamic models or conservative assumptions, fall short in capturing the complexities of real-world driving scenarios. Our methodology, which combines real-time environmental data with state-of-the-art trajectory prediction networks, offers a more effective and accurate prediction of DOS. Capable of adapting to evolving road conditions and traffic dynamics, this approach provides a comprehensive and nuanced risk assessment, which is essential for ensuring the safety of autonomous driving systems.
@@
  • We propose a novel approach to predict DOS, anchored in advanced trajectory prediction networks, and coupled with an adaptable representation of DOS. This method effectively expands the predicted trajectories into a DOS, maintaining the fidelity of the trajectory predictions.
  • We establish defined objectives for the design of DOS, focusing on coverage and area, which are instrumental in influencing the safety and efficiency of autonomous driving. A specialized training strategy has been developed in line with these objectives, enabling simultaneous optimization of both safety and efficiency.
  • We conducted a series of qualitative and quantitative experiments to validate the superiority of the proposed DOS and explore the characteristics and upper limits of different DOS representations.
@@ II-A Trajectory Prediction Networks
In the realm of autonomous driving systems, the accuracy of trajectory prediction networks is paramount in ensuring traffic safety [1, 2]. These networks utilize deep learning techniques to process complex time-series data and predict the trajectories of vehicles and pedestrians across different time scales. Typical networks, such as Recurrent Neural Networks (RNNs) [3, 4] and Convolutional Neural Networks (CNNs) [5, 6], are employed to enhance the effectiveness of trajectory prediction by handling temporal sequences and spatial features. Furthermore, by constructing social models or interaction models [7, 8], the interactions between individuals, such as pedestrian avoidance and coordinated vehicle movement, can be taken into account. Additionally, through the incorporation of attention mechanisms and graph neural networks [9, 10], these trajectory prediction networks exhibit higher accuracy and robustness in handling complex interaction dynamics in dense traffic environments.
However, relying solely on predicted trajectory information falls short, as trajectory prediction results often contain inherent errors, especially in challenging or unknown scenarios where predictions may exhibit significant inaccuracies [11]. To address this challenge, our study introduces DOS prediction to rectify potentially inaccurate trajectory predictions. This enhancement bolsters the autonomous driving system’s capability to contend with uncertainty in complex environments, providing comprehensive information support for vehicle planning and decision-making.
@@ II-B Online Estimation for Trajectory Prediction Performance
In the field of autonomous driving, recent scholarly research has underscored the significance of online monitoring algorithms in addressing predictive deficiencies in trajectory prediction models. These algorithms encompass uncertainty-based approaches [12, 13], anomaly detection techniques [14, 15], and error estimation methods [16]. Uncertainty-based algorithms [17] aim to quantify prediction uncertainty using probabilistic models, enabling autonomous vehicles to assess the reliability of trajectory predictions and make informed decisions in uncertain environments. Anomaly detection techniques [18] focus on identifying deviations from in-distribution domain, promptly flagging potential anomaly or prediction failures for corrective actions. Error estimation approaches [19] calculate prediction error margins, providing insights into deviation extents and facilitating contingency planning. Together, these monitoring approaches enhance the robustness and safety of autonomous driving systems, enabling them to adapt effectively to unpredictable traffic situations.
Nonetheless, the existing methods lack exploration of how to rectify trajectory prediction results based on estimation outcomes. To address this challenge, this study introduces a novel method for predicting DOS, offering an effective means to address the limitations of trajectory prediction models. This method, without compromising the prediction accuracy of existing trajectory prediction networks, transforms predicted trajectories into a more reliable and trustworthy form.
@@ II-C Set-based Prediction
The prediction of occupancy sets plays a pivotal role in analyzing the safety of autonomous driving systems. Traditional occupancy set prediction methods are often based on physical models and dynamic system analysis, suitable for static and predictable environments. For instance, by combining formalized traffic rules and non-deterministic motion models, these methods predict the positions and velocity sets that vehicles, pedestrians, and bicycles might occupy [20, 21, 22]. As technology advances, new research directions involve predicting occupancy sets in more dynamic and uncertain environments, especially by combining probabilistic models and machine learning methods to enhance adaptability and accuracy [23, 24, 25]. These modern approaches are better equipped to handle environmental uncertainty and traffic complexity, particularly in complex scenarios involving multiple vehicles and pedestrians. Probabilistic prediction models offer a probabilistic representation of the predicted occupancy set, but they may not provide reasonable information about the extent of the occupancy set.
This study introduces a learning-based DOS prediction method, which is more flexible and efficient compared to traditional set-based prediction methods. Furthermore, traditional occupancy set evaluations often lack comprehensive and well-defined assessment metrics. To address this issue, this research proposes two new evaluation metrics, coverage and occupancy set area, providing clear evaluation references for this field of study. Additionally, to achieve the aforementioned objectives, a multi-objective optimization training strategy is introduced, ensuring both the safety and efficiency of autonomous driving.
@@ III-A Trajectory Prediction
Figure 2: Trajectory prediction network with DOS prediction.
Our research builds upon the rapidly evolving field of trajectory prediction, leveraging its advanced techniques while addressing the inevitable issue of prediction errors. Consequently, we provide an estimation of a trajectory-prediction-aware DOS.
In a representative trajectory prediction network $F$, the process begins with meticulous input data preprocessing, primarily focusing on historical location information $S={\textbf{s}_{i}^{-T_{h}:0}}$, including coordinates, velocity, and direction of the $i$-th traffic participant at time $t$, where $i$ ranges from 1 to $N$. This preprocessing phase encompasses feature extraction, affine transformations of trajectory data, processing of local map data $\mathcal{M}$, and masking of incomplete historical or future trajectories. It plays a critical role in structuring the input data for effective analysis. The prediction results $\hat{Y}$ include the predicted trajectories $\hat{Y}_{i}=\hat{\textbf{y}}_{i}^{1:T_{f}}$ for traffic participant $A_{i}$, which are determined by $\hat{Y}=F(S,\mathcal{M})$.
@@
The effectiveness of the trajectory prediction model is contingent upon its optimization and training process. It undergoes rigorous training using extensive datasets that encompass a wide spectrum of movement patterns, enabling it to learn and adapt to various behaviors. Key training techniques, including backpropagation and gradient descent, are employed to refine network weights. Hyperparameter tuning, involving adjustments to learning rates, batch sizes, and network layers, is meticulously conducted to optimize performance.
@@ III-B DOS Representation
The center $\mu_{i}^{t}$ represents the expected position of the traffic participant at time $t$, which is predicted based on the trajectory prediction model. The major and minor axis lengths, $l_{i}^{t}$ and $w_{i}^{t}$, determine the size of the ellipse, representing the uncertainty in both longitudinal and lateral directions. The orientation angle $\theta_{i}^{t}$ captures the direction of this uncertainty.
In the design of DOS, different assumptions can be applied to impose varying levels of constraints:
@@
These mathematical representations of DOS allow us to effectively capture and manage uncertainty in the context of traffic participant movements, providing a structured and interpretable framework for safety-aware autonomous systems.
@@ III-C DOS Prediction Module
The core of our approach lies in the design of a DOS prediction module, which leverages information from the trajectory prediction network. This module is responsible for generating DOS for traffic participant within the prediction horizon $T_{f}$. In this subsection, we delve into the details of our model design.
Our model architecture is based on neural networks, which have demonstrated remarkable effectiveness in various domains for capturing complex relationships. For each traffic participant, our model’s input includes the output features $H$ from the encoding layer of the trajectory prediction network and the predicted trajectory $\hat{Y}_{i}$. $H$ aggregates the encoding of all input information from the trajectory prediction network, such as modeling of maps, interactions, and historical data, thus providing crucial references for DOS prediction. $\hat{Y}_{i}$ serves as supplementary information.
Given the advantages of GRUs in capturing temporal dependencies and sequence patterns, we employ GRUs as the primary structure in the DOS prediction module. As illustrated in Fig. 3, $H$ is first duplicated along the time dimension to serve as input to the GRU-based encoder, resulting in encoded features $H_{1}$. On the other hand, $Y_{i}$ goes through a linear layer and a GRU-based encoder to obtain features $H_{2}$. Subsequently, $H_{1}$ and $H_{2}$ are temporally concatenated as input to the decoder for DOS prediction. Furthermore, specific activation functions are designed to ensure that the predicted ellipse parameters, including axis lengths and orientation angles, fall within reasonable ranges.
@@ III-D Design Objectives
To achieve these design objectives, we have designed specific loss functions to separately consider coverage and area:
@@
$L_{c}$ is formulated based on coverage of the DOS. For each time step and each predicted object, we assess coverage by computing the Mahalanobis distance $d_{M}$ between the true position $\textbf{p}_{i}^{t}$ and the predicted DOS. By encouraging $d_{M}$ to decrease, we aim to boost coverage, especially when $d_{M}>1$, indicating that the DOS does not envelop $\textbf{p}_{i}^{t}$, and in such cases, it should be penalized more rigorously. Specifically, for a given prediction:
${l}_{i}^{t}=\begin{cases}{d}_{M}&\text{if }{d}_{M}\leq 1\\ {{d}_{M}}^{\alpha}+\beta&\text{if }{d}_{M}>1\end{cases},$ (2)
where $\alpha$ and $\beta$ represent factors that can be manually set or obtained through data-driven statistics. ${L}_{c}=\frac{1}{NT}{\sum}_{i=1}^{N}{\sum}_{t=1}^{T}{{l}_{i}^{t}}$.
@@ IV-B Evaluation Metrics
    $CR_{t}=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}1,&\text{if }\textbf{p}_{i}^{t}\text{ is inside DOS at }t\\ 0,&\text{otherwise}\end{cases}$ (4)
@@
CR OSA
1s
2s
3s
4s
1s
2s
3s
4s
PR-based ($z=0.5$)
0.417
0.352
0.327
0.254
0.179
1.109
2.667
5.309
PR-based ($z=1.0$)
0.762
0.725
0.716
0.627
0.715
4.434
10.666
21.237
PR-based ($z=1.5$)
0.899
0.877
0.883
0.829
1.608
9.976
23.999
47.784
PR-based ($z=2.0$)
0.951
0.938
0.945
0.919
2.859
17.737
42.664
84.949
SA-based ($z=1.0$)
0.460
0.428
0.445
0.401
0.319
2.146
6.055
13.118
SA-based ($z=1.5$)
0.654
0.623
0.637
0.595
0.717
4.827
13.623
29.515
SA-based ($z=2.0$)
0.778
0.754
0.769
0.733
1.274
8.582
24.219
52.471
Circle-based ($r=1.0$)
0.917
0.753
0.695
0.628
3.142
3.142
3.142
3.142
Circle-based ($r=1.5$)
0.963
0.824
0.754
0.684
7.069
7.069
7.069
7.069
Circle-based ($r=2.0$)
0.981
0.873
0.800
0.721
12.566
12.566
12.566
12.566
Circle-v1-based ($z=0.5$)
0.792
0.753
0.754
0.721
0.785
3.142
7.069
12.566
Circle-v1-based ($z=0.7$)
0.858
0.811
0.808
0.769
1.539
6.158
13.854
24.63
Circle-v1-based ($z=1.0$)
0.917
0.873
0.865
0.826
3.142
12.566
28.274
50.265
Circle-v2-based ($z=1.0$)
0.708
0.719
0.731
0.73
0.350
2.107
5.141
14.234
Circle-v2-based ($z=1.5$)
0.792
0.788
0.793
0.788
0.787
4.740
11.567
32.026
DOS (circle)
0.960
0.882
0.853
0.787
3.369
7.815
12.633
18.849
DOS (elipse, $\theta=0$)
0.966
0.912
0.880
0.810
2.303
5.676
9.004
13.087
DOS (elipse, $\theta=\pi/4$)
0.958
0.877
0.846
0.781
3.442
7.591
12.033
17.844
DOS (elipse, v1)
0.967
0.904
0.87
0.800
2.869
6.191
9.483
13.813
DOS (elipse, v2)
0.968
0.909
0.877
0.807
2.453
5.706
9.088
13.448
DOS (elipse)
0.969
0.912
0.878
0.811
2.461
5.827
9.147
13.364
@@
CR OSA
1s
4s
1s
4s
PR-based ($z=1.0$)
0.408
0.25
0.46
9.637
Circle-based ($r=1.0$)
0.937
0.604
3.142
3.142
Circle-v1-based ($z=1.0$)
0.937
0.869
3.142
50.265
Circle-v2-based ($z=1.5$)
0.830
0.793
0.872
21.051
DOS (elipse, $\theta=0$)
0.958
0.791
1.590
9.885
DOS (elipse, $\theta=\pi/4$)
0.942
0.778
2.106
11.542
DOS (elipse, v1)
0.940
0.779
2.094
11.088
DOS (elipse, v2)
0.953
0.794
1.874
10.421
DOS (elipse)
0.956
0.786
1.617
9.751

### v2

@@ Abstract
In the rapidly evolving field of autonomous driving, reliable prediction is pivotal for vehicular safety. However, trajectory predictions often deviate from actual paths, particularly in complex and challenging environments, leading to significant errors. To address this issue, our study introduces a novel method for Dynamic Occupancy Set (DOS) prediction, it effectively combines advanced trajectory prediction networks with a DOS prediction module, overcoming the shortcomings of existing models. It provides a comprehensive and adaptable framework for predicting the potential occupancy sets of traffic participants. The innovative contributions of this study include the development of a novel DOS prediction model specifically tailored for navigating complex scenarios, the introduction of precise DOS mathematical representations, and the formulation of optimized loss functions that collectively advance the safety and efficiency of autonomous systems. Through rigorous validation, our method demonstrates marked improvements over traditional models, establishing a new benchmark for safety and operational efficiency in intelligent transportation systems.
@@ I Introduction
The rapid progression of autonomous driving technology is heralding a new era in transportation systems, bringing forth notable enhancements in traffic efficiency and vehicular safety. Yet, ensuring the safety and reliability of autonomous vehicles in the multifaceted real-world scenarios presents a significant challenge. This complexity arises not only from the advanced nature of the technology but also from the necessity to sustain system reliability in intricate environments. The realm of trajectory prediction networks has undergone substantial development, becoming fundamental in boosting the predictive capabilities essential for autonomous vehicle navigation. These networks employ advanced algorithms and extensive datasets, enabling predictions of the movements of traffic participants by considering various factors like vehicle dynamics, road conditions, and traffic patterns. Enhanced prediction accuracy and adaptability of these networks are particularly valuable in complex urban environments or situations involving unpredictable human interactions. However, the complexity of real-world conditions and inherent limitations of predictive models [1] often lead to discrepancies between predicted and actual trajectories, posing a detriment to the reliable assurance of autonomous driving safety.
@@
In response to above challenges, our study introduces an novel Dynamic Occupancy Set (DOS) representation and estimation method for traffic participants. DOS, which defines all potential positions a vehicle may occupy within a specified timeframe, plays a pivotal role in reinforcing the safety framework of autonomous driving systems. Traditional set-based prediction methods, often reliant on simplistic dynamic models or conservative assumptions, fall short in capturing the complexities of real-world driving scenarios. Our methodology, which integrates real-time environmental data with advanced trajectory prediction networks, enhances DOS predictions by minimizing the DOS area while maximizing coverage of traffic participants’ future locations, thus balancing efficiency and safety, as shown in Fig. 1. Capable of adapting to evolving road conditions and traffic dynamics, this approach provides a comprehensive and nuanced risk assessment, which is essential for ensuring the safety of autonomous driving systems.
@@
  • Integration of DOS Prediction in Trajectory Prediction Networks. This study pioneers the integration of DOS prediction with advanced trajectory prediction networks, enhancing the robustness and accuracy of autonomous vehicle navigation under uncertain conditions.
  • Mathematical Model of DOS. The research introduces an innovative elliptic set representation of DOS, enabling precise and flexible depiction of spatial uncertainties in traffic participant movements.
  • Optimized Objectives for DOS Prediction. The paper sets clear design objectives for DOS, emphasizing coverage and area to enhance the safety and efficiency of autonomous driving. A specialized training strategy aligns with these goals, ensuring optimal balance between safety and operational efficiency.
@@ II-A Trajectory Prediction Networks
In autonomous driving, the accuracy of trajectory prediction networks is crucial for ensuring traffic safety [2, 3]. These networks use deep learning to process complex time-series data, predicting vehicle and pedestrian movements over various time scales. Recurrent Neural Networks (RNNs) [4] and Convolutional Neural Networks (CNNs) [5] manage temporal sequences and spatial features effectively. Additionally, social and interaction models [6, 7] capture dynamics like pedestrian avoidance and vehicle coordination. The integration of attention mechanisms and graph neural networks [8, 9] further enhances the accuracy and robustness in dense traffic scenarios. Despite these advances, trajectory predictions often contain errors, particularly in complex or unfamiliar environments. To mitigate this, our study introduces DOS prediction to improve reliability in vehicle planning and decision-making.
@@ II-B Online Estimation for Trajectory Prediction Performance
Recent research highlights the importance of online monitoring algorithms to address deficiencies in trajectory prediction models. These methods include uncertainty-based approaches [10, 11], anomaly detection techniques [12], and error estimation methods [13], which help quantify prediction uncertainty [14], detect potential anomalies or outliers [15], and calculate prediction error margins [16], respectively. These tools improve the robustness and adaptability of autonomous driving systems to unpredictable traffic conditions. Nonetheless, existing methods have not thoroughly explored how to compensate for deficiencies in trajectory prediction results based on estimation outcomes. To address this challenge, this study introduces a novel method for predicting DOS, providing an effective means to overcome the limitations of trajectory prediction models. This method preserves the prediction accuracy of existing trajectory prediction networks while transforming predicted trajectories into a more reliable and trustworthy form.
@@ II-C Set-based Prediction
The prediction of occupancy sets is fundamental in assessing autonomous driving safety. Traditionally, these predictions rely on physical models and formalized traffic rules [17, 18, 19], suitable for static and predictable scenarios. With advances in technology, new approaches using probabilistic models and machine learning have emerged to predict occupancy sets in more dynamic and uncertain contexts [20, 21]. These methods provide a more adaptable representation of potential occupancies but may lack precise information about the extent of these sets. Additionally, self-aware prediction model [22] estimates prediction errors to monitor performance and define error bounds. Overall, probabilistic prediction models that aim for maximum likelihood estimation (MLE) and error estimation models guided by actual prediction errors do not guarantee maximum coverage of actual future trajectories within the smallest possible space, which is critical for constructing reliable DOS.
Our research introduces a novel learning-based DOS prediction method, which is more flexible and effective than traditional methods. We also propose new evaluation metrics—coverage and occupancy set area—to provide clearer assessment criteria. A multi-objective optimization training strategy is implemented to enhance both the safety and efficiency of autonomous driving.
@@ III-A Trajectory Prediction
Figure 2: Trajectory-prediction-aware DOS prediction model.
Our research builds upon the rapidly evolving field of trajectory prediction, leveraging its advanced techniques while addressing the inevitable issue of prediction errors. Consequently, we propose a trajectory-prediction-aware DOS prediction model, as shown in Fig. 2.
In a representative trajectory prediction network $F$, the process begins with meticulous input data preprocessing, primarily focusing on historical location information $\textbf{S}={\textbf{s}_{i}^{-T_{h}:0}}$, including coordinates, velocity, and direction of the $i$-th traffic participant at time $t$, where $i$ ranges from 1 to $N$. This preprocessing phase encompasses feature extraction, affine transformations of trajectory data, processing of local map data $\mathcal{M}$, and masking of incomplete historical or future trajectories. It plays a critical role in structuring the input data for effective analysis. The prediction results $\hat{\textbf{Y}}$ include the predicted trajectories $\hat{\textbf{Y}}_{i}=\hat{\textbf{y}}_{i}^{1:T_{f}}$ for traffic participant $A_{i}$, which are determined by $\hat{\textbf{Y}}=F(S,\mathcal{M})$.
@@
The trajectory prediction model undergoes rigorous training using extensive datasets that encompass a wide spectrum of movement patterns, enabling it to learn and adapt to various behaviors. Key training techniques, including backpropagation and gradient descent, are employed to refine network weights. Hyperparameter tuning, involving adjustments to learning rates, batch sizes, and network layers, is meticulously conducted to optimize performance.
@@ III-B DOS Representation
The ellipse center $\mu_{i}^{t}$ represents the most probable position of the traffic participant at time $t$, which is extracted from the predicted trajectory. The major and minor axis lengths, $l_{i}^{t}$ and $w_{i}^{t}$, determine the size of the ellipse, representing the uncertainty in both longitudinal and lateral directions. The orientation angle $\theta_{i}^{t}$ captures the direction of this uncertainty. Different constraints can be applied to achieve diverse variants of DOS based on varying assumptions:
@@
These DOS representations allow us to effectively capture and manage uncertainty in the context of traffic participant movements, providing a structured and interpretable framework for safety-aware autonomous systems.
@@ III-C DOS Prediction Module
The core of our approach is centered on the development of a DOS prediction module that utilizes information from the trajectory prediction network. This module is tasked with generating the DOS for traffic participants across the prediction horizon $T_{f}$, represented as $\mathcal{O}_{1:N}^{1:T_{f}}.$
Our model architecture is based on neural networks, which have demonstrated remarkable effectiveness in various domains for capturing complex relationships. For each traffic participant, our model’s input includes the output features H from the encoding layer of the trajectory prediction network and the predicted trajectory $\hat{\textbf{Y}}_{i}$. H aggregates the encoding of all input information from the trajectory prediction network, such as modeling of maps, interactions, and historical data, thus providing crucial references for DOS prediction. $\hat{\textbf{Y}}_{i}$ serves as supplementary information.
Given the advantages of GRUs in capturing temporal dependencies and sequence patterns, we employ GRUs as the primary structure in the DOS prediction module. As illustrated in Fig. 3, H is first duplicated along the time dimension to serve as input to the GRU-based encoder, resulting in encoded features $\textbf{H}_{1}$. On the other hand, $Y_{i}$ goes through a linear layer and a GRU-based encoder to obtain features $\textbf{H}_{2}$. Subsequently, $\textbf{H}_{1}$ and $\textbf{H}_{2}$ are temporally concatenated as input to the decoder for DOS prediction. Furthermore, specific activation functions are designed to ensure that the predicted ellipse parameters, including axis lengths and orientation angles, fall within reasonable ranges.
@@ III-D Design Objectives
To achieve the above objectives, we design specific loss functions to separately consider coverage and area:
@@
$L_{c}$ is formulated based on the coverage of DOS. For each time step and each predicted object, we construct a multivariate gaussian distribution based on the predicted DOS, with its mean and covariance determined by the center and shape of $\mathcal{O}_{i}^{t}$. Coverage is then assessed by calculating the Mahalanobis distance $d_{M}$ between the true position $\textbf{p}_{i}^{t}$ and the center of the distribution. By encouraging $d_{M}$ to decrease, we aim to boost coverage, especially when $d_{M}>1$, indicating that the DOS does not envelop $\textbf{p}_{i}^{t}$, and in such cases, it should be penalized more rigorously. Specifically:
${L}_{c}=\frac{1}{NT}{\sum}_{i=1}^{N}{\sum}_{t=1}^{T}{{l}_{c,i}^{t}},$ (2)
${l}_{c,i}^{t}=\begin{cases}{d}_{M}&{d}_{M}\leq 1\\ {{d}_{M}}^{\alpha}+\beta&{d}_{M}>1\end{cases},$ (3)
where $\alpha$ and $\beta$ represent factors that can be manually set or obtained through data-driven statistics.
@@ IV-B Evaluation Metrics
    $CR_{t}=CR_{i}^{t}$ (5)
    $CR_{i}^{t}=\begin{cases}1&\text{if }\textbf{p}_{i}^{t}\text{ is inside DOS at }t\\ 0&\text{otherwise}\end{cases}$ (6)
@@
CR (%) OSA
1s 2s 3s 4s 1s 2s 3s 4s
PR-based ($z=0.5$) 41.7 35.2 32.7 25.4 0.179 1.109 2.667 5.309
PR-based ($z=1.0$) 76.2 72.5 71.6 62.7 0.715 4.434 10.666 21.237
PR-based ($z=1.5$) 89.9 87.7 88.3 82.9 1.608 9.976 23.999 47.784
PR-based ($z=2.0$) 95.1 93.8 94.5 91.9 2.859 17.737 42.664 84.949
SA-based ($z=1.0$) 46.0 42.8 44.5 40.1 0.319 2.146 6.055 13.118
SA-based ($z=1.5$) 65.4 62.3 63.7 59.5 0.717 4.827 13.623 29.515
SA-based ($z=2.0$) 77.8 75.4 76.9 73.3 1.274 8.582 24.219 52.471
Circle-based ($r=1.0$) 91.7 75.3 69.5 62.8 3.142 3.142 3.142 3.142
Circle-based ($r=2.0$) 98.1 87.3 80.0 72.1 12.566 12.566 12.566 12.566
Circle-v1-based ($z=0.7$) 85.8 81.1 80.8 76.9 1.539 6.158 13.854 24.63
Circle-v1-based ($z=1.0$) 91.7 87.3 86.5 82.6 3.142 12.566 28.274 50.265
Circle-v2-based ($z=1.0$) 70.8 71.9 73.1 73.0 0.350 2.107 5.141 14.234
Circle-v2-based ($z=1.5$) 79.2 78.8 79.3 78.8 0.787 4.740 11.567 32.026
DOS (circle) 96.0 88.2 85.3 78.7 3.369 7.815 12.633 18.849
DOS (elipse, $\theta=0$) 96.6 91.2 88.0 81.0 2.303 5.676 9.004 13.087
DOS (elipse, $\theta=\pi/4$) 95.8 87.7 84.6 78.1 3.442 7.591 12.033 17.844
DOS (elipse, v1) 96.7 90.4 87.0 80.0 2.869 6.191 9.483 13.813
DOS (elipse, v2) 96.8 90.9 87.7 80.7 2.453 5.706 9.088 13.448
DOS (elipse) 96.9 91.2 87.8 81.1 2.461 5.827 9.147 13.364
@@
CR(%) OSA
1s 4s 1s 4s
PR-based ($z=2.0$) 63.8 45.6 1.841 38.55
SA-based ($z=1.0$) 94.1 77.1 1.956 11.469
Circle-based ($r=1.0$) 93.7 60.4 3.142 3.142
Circle-v1-based ($z=1.0$) 93.7 86.9 3.142 50.265
Circle-v2-based ($z=1.5$) 83.0 79.3 0.872 21.051
DOS (elipse, $\theta=0$) 95.8 79.1 1.590 9.885
DOS (elipse, $\theta=\pi/4$) 94.2 77.8 2.106 11.542
DOS (elipse, v1) 94.0 77.9 2.094 11.088
DOS (elipse, v2) 95.3 79.4 1.874 10.421
DOS (elipse) 95.6 78.6 1.617 9.751